OpenStudy (anonymous):

Write an equation for the line tangent to the circle x2 + y2 = 17, at the point (4, 1). Enter your answer in point-slope form: y - y1 = m(x - x1); example: y+3=-5(x-2)

OpenStudy (amistre64):

y - 1 = f'(4)(x-4)

OpenStudy (anonymous):

y1=1 x1=4 m=dy/dx at (4,1)

OpenStudy (amistre64):

f'(x) creates 2 points of tangency to a circle

OpenStudy (amistre64):

Dx{x^2 + y^2 = 17} = 2x + 2y.y' = 0 y' = -x/y ; at (4,1) .... ok, I forgot to read that single point part lol. y-1 = -4(x-4)

OpenStudy (anonymous):

i just did this prob!!